Two metal balls A and B of negligible radius are floating at rest on Space Station Freedom between two metal bulkheads, connected by a taut nonconducting thread of length 2.70 m. Ball A carries charge q, and ball B carries charge 2q. Each ball is 1.16 m away from a bulkhead.

(a) If the tension in the string is 2.05 N, what is the magnitude of q? (Ignore any contribution from induced charge in the bulkheads.)

Answer in microC
Answer #1

The tension within the string is the same as the force felt between the two objects. The force felt between two point objects due to there electrostatic repulsion is given by

F = kq1q2/d^2

2.05 = 9*10^9*q*2q/2.7^2

q^2 = 8.3*10^(-10)

q = 2.88*10^(-5) = 28.8 microC